Solve the quadratic equation by completing the square:
2x^2 + 10x + 11 = 0
Can someone please help me out on this, thanks a lot!

Answer by uma(370)

Given 2x^2 + 10x + 11 = 0
==> x^2 + 5x = -11/2 [dividing by 2 throughout]
==> x^2 + 5x + (5/2)^2 = -11/2 + (5/2)^2 [adding (5/2)^2 to complete the square]
==> (x+ 5/2)^2 = -11/2 + 25/4
==> (x + 5/2)^2 = (-22 + 25)/4 [taking LCD]
==> (x + 5/2)^2 = 3/4
==> x + 5/2 = +/- sqrt(3)/2 [taking sq root]
==> x = - 5/2 +/- sqrt(3)/2
Good Luck!!!